# Diffusion Implicit Policy for Unpaired Scene-aware Motion Synthesis

### v1

@@ Abstract
Human motion generation is a long-standing problem, and scene-aware motion synthesis has been widely researched recently due to its numerous applications. Prevailing methods rely heavily on paired motion-scene data whose quantity is limited. Meanwhile, it is difficult to generalize to diverse scenes when trained only on a few specific ones. Thus, we propose a unified framework, termed Diffusion Implicit Policy (DIP), for scene-aware motion synthesis, where paired motion-scene data are no longer necessary. In this framework, we disentangle human-scene interaction from motion synthesis during training and then introduce an interaction-based implicit policy into motion diffusion during inference. Synthesized motion can be derived through iterative diffusion denoising and implicit policy optimization, thus motion naturalness and interaction plausibility can be maintained simultaneously. The proposed implicit policy optimizes the intermediate noised motion in a GAN Inversion manner to maintain motion continuity and control keyframe poses though the ControlNet branch and motion inpainting. For long-term motion synthesis, we introduce motion blending for stable transitions between multiple sub-tasks, where motions are fused in rotation power space and translation linear space. The proposed method is evaluated on synthesized scenes with ShapeNet furniture, and real scenes from PROX and Replica. Results show that our framework presents better motion naturalness and interaction plausibility than cutting-edge methods. This also indicates the feasibility of utilizing the DIP for motion synthesis in more general tasks and versatile scenes. https://jingyugong.github.io/DiffusionImplicitPolicy/
footnotetext: Corresponding Author.
1 Introduction
@@
Figure 1: Policy learning frameworks. (a) Explicit policy is trained with paired motion-scene data given task to predict the final motion. (b) Implicit policy optimizes the motion from initialization accordingly. (c) Diffusion policy gradually denoise the motion based on current scene, task, and noised motion. (d) Our diffusion implicit policy iteratively denoises and optimizes the motion to ensure motion naturalness, diversity, interaction plausibility simultaneously without need for any paired motion-scene data.
@@
Synthesizing human motion in real 3D scenes has attracted significant attention in recent years [4, 42, 43, 62], due to its wide applications in scene simulation, digital human animation, and virtual/augmented reality.
@@
Thanks to learning-based 3D perception [32, 60], pioneers [4, 42, 62] have attempted to synthesize motion in scenes with feasible human-scene interaction. However, in almost all previous works, paired motion-scene data are required to learn scene-aware motion policies. The majority of prevailing methods [37, 62] learn Explicit Policies to directly predict the desired motion based on current states and goals (Fig. 1 (a)). Some of them [42, 43] utilized a second-stage Implicit Policy optimization but sacrifice the motion naturalness for interaction plausibility (Fig. 1 (b)). Recent works [18, 45] utilized conditional Diffusion Policies to achieve better performance (Fig. 1 (c)), where massive paired motion-scene data is also necessary.
@@
In fact, captured human motion data [24, 22] is far more abundant than paired motion-scene data [13, 44]. Motion synthesis that relies heavily on paired data will inevitably suffer from limited diversity. Meanwhile, the generalization ability is hard to guarantee when trained on limited scenes and applied to various other scenes.
@@
Based on this observation, we propose a unified framework, termed Diffusion Implicit Policy (DIP) (Fig. 1 (d)), which disentangles human-scene interaction from motion synthesis during training and then integrate motion denoising with implicit policy optimization during inference. In this way, paired motion-scene data is no longer necessary for training, and motion naturalness and interaction plausiblity can be ensured simultaneously for scene-aware motion synthesis.
@@
In the DIP, a motion diffusion model is employed to make the synthesized motion more and more natural throughout the entire denoising process(Fig. 2 (b)). We equipped the diffusion model with a ControlNet [51] branch to provide keyframe joint hints for historical motion and future goals. Following previous works [39, 46], the diffusion model is designed to predict the original motion at each denoising step and then sample the denoised motion from a normal distribution accordingly. Thus, we can well utilize the stochastic process to pursue plausible human-scene interactions. Specifically, interaction-based reward functions are designed to assess the consistency between motions and scenes. These reward functions are used as implicit policy to optimize the sampling distribution, ensuring that the sampled denoised motion corresponds better to the 3D scenes at each denoising step. As the denoising process can also be viewed as optimization for motion naturalness, the entire scene-aware motion synthesis can be framed as an optimization problem to simultaneously pursue both motion naturalness and interaction plausibility.
@@
In addition, for long-term motion synthesis involving multiple tasks, we need to take historical motion as constrain when synthesizing future motion. To maintain continuity between historical and future motions, we employ a time-variant motion blending, where we interpolate the rotation matrix in the power space and the translation in standard linear space. Thus far, the proposed framework can synthesize long-term motion in general scenes without any training on paired motion-scene data.
For performance evaluation, we use scenes cluttered with furniture from ShapeNet [6] to assess the ability on human-object interaction. We also take PROX [13] and Replica [38] to demonstrate the generalization ability in scene-aware motion synthesis. We compared the proposed method with prevailing works based on physical and perceptual scores. Comprehensive experiments support our claims and indicate that the synthesized motion produced by the proposed method demonstrates better performance.
@@
2 Related Work
[…]
@@ 2 Related Work
Motion Synthesis. Motion synthesis is a long-standing problem that has been studied for a significant period [8, 17, 37]. This topic has been researched conditioned on various signals, including motion prefixes [25], actions [11, 29, 48], music [10, 41], and text [30, 12, 39]. Action2Motion [11] employed a recurrent conditional VAE for motion creation, where historical data was utilized to predict the subsequent pose. ACTOR [29] encoded the entire motion sequence into a latent feature space, significantly reducing the accumulative error in recurrent methods. TEMOS [30] utilized a VAE to learn a shared latent space for motion and textual description. The motion distribution and text distribution were well aligned by minimizing the KL divergence. T2M [12] further trained a text-to-length estimator, enabling the network to automatically predict the length of the generated motion. MDM [39] and MotionDiffuse [54] are the pioneers to leverage diffusion model for human motion synthesis. Subsequent works [7, 47, 52, 9, 53, 20, 46] further improved the controllability and quality of the generated results through database retrieval, spatial control, fine-grained captioning.
Thanks to the advancements in motion synthesis, we follow the MDM [39] and extend it to scene-aware motion synthesis, using interaction-based reward functions as an implicit policy.
[…]
@@
3 Method
3.1 Preliminary
Motion Representation. For human motion, we take the SMPL-X model [28] to represent the pose at each frame. Here, we mainly consider the global orientation represented in axis-angle $\theta_{global}\in\mathbb{R}^{3}$, joint rotation in axis-angle $\theta_{j=1:21}\in\mathbb{R}^{63}$ and the translation $\tau\in\mathbb{R}^{3}$. Accordingly, for each frame $s$, the human pose can be defined as $P_{s}=\{\theta_{s,global},\theta_{s,j=1:21},\tau_{s}\}\in\mathbb{R}^{69}$, and the synthesized motion consisting of consecutive poses can be annotated as $\hat{P}=\{\hat{P}_{s}\}_{s=1:S}$. The body shape $\beta\in\mathbb{R}^{10}$ and hand pose $\theta_{h}\in\mathbb{R}^{24}$ are always keep the same as initial human body for simplicity. The first $K$ joints $J=J_{1:K}\in\mathbb{R}^{K\times 3}$ and body mesh with $V$ vertices $M(\tau,\theta_{global},\beta,\theta_{j},\theta_{h})\in\mathbb{R}^{V\times 3}$ are taken as auxiliary representation for human pose.
@@ 3.1 Preliminary
where $\mathcal{N}$ is a normal distribution and $\alpha_{t=1:T}$ are a series of hyper-parameters. The distribution of final noised motion $x_{T}$ will approximate to $\mathcal{N}(0,I)$. Like MDM [39], we train a diffusion model to predict the original motion directly
@@
where $t$ is the time step and $a$ is the action label for easier control of human motion behavior. As for motion synthesis, the denoising procedure can be formulated as:
$P(x_{t-1}|\hat{x}_{0}^{\phi},x_{t})=\mathcal{N}(x_{t-1};\mu_{t}(\hat{x}_{0}^{\phi},x_{t}),\tilde{\beta}_{t}\mathbf{I}),$ (3)
@@
3.2 Overview
Figure 2: (a) indicates the overall pipeline of scene-aware motion synthesis. Any feasible command will be first decomposed into sub-task with action-object pairs. Then, we will synthesize the future motion according to historical motion and current sub-task. Last, the synthesized motions will be fused into the historical motion to obtain the final long-term motion. (b) presents the framework of Diffusion Implicit Policy (DIP). In each iteration of the DIP, the diffusion model will denoise the motion and enable the synthesized motion to appear more natural, and implicit policy optimization from reward will endow the motion with plausible interaction. The random sampling step can help the framework synthesize motion with diverse styles.
@@ 3.2 Overview
We can first decompose the whole command into a list of sub-task (interaction behavior and object pair) to be finished via current LLMs [1, 40]. For each sub-task, the human may go to some place or interact with the objects in the scene (e.g., sitting on the chair).
@@
Given a sub-task, we will first locate the goal position as COINS [61]. Then, we will judge current action, and fetch reward functions for implicit policy accordingly.
@@
We train a diffusion model conditioned on motion action and keyframe joints to synthesize natural motion with easy control, where body meshes and joints are fully differentiable for Diffusion Implicit Policy (Sec. 3.3). Later, we model the interaction-based reward functions (Sec. 3.4) and take them to optimize the sampling distribution of denoised motion at each denoising step (Sec. 3.5). The motion prior from diffusion model and implicit policy from reward function are integrated together to synthesize scene-aware motion with desired interactions (as shown in Fig. 2 (b)).
@@
Given historical motion, the synthesized future motion should be consistent with it. Thus, we design to derive the long-term motion via a time-variant blending (Sec. 3.6) where translation are interpolated linearly and rotation are blended in the matrix power space for motion transition. By now, we can synthesize long-term motion in 3D scenes when all sub-tasks are completed.
3.3 Conditional Diffusion Model
@@
To simplify the formulation of reward functions in Diffusion Implicit Policy, we train diffusion model $\phi$ (Fig. 3 (a)) to predict the original motion $\hat{x}_{0}^{\phi}$ for each noised $x_{t}$ following MDM [39] as shown in Eq. 2. We also take $x_{0}\in\mathbb{R}^{S\times 69}$ consisting of human translation, orientation, and joint rotations in $S$ frames as the representation of motion.
@@
For any original motion $x_{0}$, we take the local coordinate of motion to reduce representation redundancy. Specifically, we translate and horizontally rotate the motion based on the human pose in the first frame. In the transformed motion, the orientation of the human body in the first frame lie in the $yz$-plane ($y\geq 0$), and the pelvis is positioned at the origin (details are available in Supplementary Material). After motion synthesis, the motion in scenes can be derived via simple translation and horizontal rotation.
@@
Figure 3: Illustration of conditional diffusion model. A diffusion model is first pre-trained conditioned on action, and then a ControlNet branch is introduced to provide keyframe joints’ hint.
@@
To better maintain consistency with historical motion and achieve future goals, we also take a ControlNet branch [51] to provide the hint of keyframe human joints (available from motion for training) that need to be controlled (Fig. 3 (b)). We choose to take the joint positions rather than joint rotation as the external hint for easier understanding of the space information in the diffusion model. Here, we mainly use the principle $22$ joints from the SMPL-X human body skeleton as the skeleton joint hint. Concretely, we will first calculate the joints’ positions according to the original human motion, and then randomly select one from these joints in a few frames as the hint and others are padded with zeros, thus the input of ControlNet branch take the form of $\mathcal{J}=\{J_{s,k}\}_{s=1:S,k=1:K}$ where non-zero values provide the controlled joint position hints.
@@
For training the ControlNet branch, all its parameters are randomly initialized, and the link layers are initialized to zero to maintain the motion synthesis capability of the main branch. Meanwhile, all parameters in original motion diffusion model $\phi$ are frozen during training. The controlled diffusion model $\varphi$ is also supervised to predict the origin motion represented by $x_{0}=\{P_{s}\}_{s=1:S}$. After fine-tuning the ControlNet branch, the controlled motion diffusion model can be formulated as
@@
For simplicity, the action $a$ and joint hint $\mathcal{J}$ are termed as the condition $c$. Later, the controlled diffusion model $\varphi$ are taken as motion prior to optimize the noised motion for higher motion naturalness.
@@
3.4 Reward Functions for Implicit Policy
Based on the aforementioned diffusion model, we introduce a set of reward functions that score the motions within scenes and serve as an implicit policy to enhance overall performance (as shown in Fig. 2 (b) blue part). Here, we primarily focus on three key aspects in the design of reward functions: motion continuity, goal achievement, and interaction plausibility.
To promote motion continuity, we introduce a Historical Motion Consistency reward $\mathcal{R}_{his}$ to ensure continuous transition between historical and synthesized motions, alongside a Smoothness reward to prevent abrupt changes with excessive acceleration $\mathcal{R}_{acc}$. Additionally, we implement a Goal Achievement reward $\mathcal{R}_{goal}$ that encourage the human to move closer to the target. We also incorporate a Contact reward $\mathcal{R}_{cont}$ and a Non-Penetration reward $\mathcal{R}_{pene}$ to facilitate plausible human-scene interactions, as well as a Non-skating reward $\mathcal{R}_{skt}$ to anchor the vertices of contact. Please kindly refer to the Supplementary Material for more details on the formulation of these reward functions.
The total reward function for implicit policy take the form of
$\begin{array}[]{rcl}\mathcal{R}_{ip}&=&\lambda_{his}\mathcal{R}_{his}+\lambda_{acc}\mathcal{R}_{acc}+\lambda_{goal}\mathcal{R}_{goal}\\ &&+\lambda_{cont}\mathcal{R}_{cont}+\lambda_{pene}\mathcal{R}_{pene}\\ &&+\lambda_{skt}\mathcal{R}_{skt}\end{array}$ (5)
@@ 3.4 Reward Functions for Implicit Policy
3.5 Diffusion Implicit Policy
As mentioned in Sec. 3.3, we utilize a diffusion model with ControlNet branch $\varphi$ to predict the origin motion $\hat{x}_{0}^{\varphi}$. Thus, similar to Eq. 3, we can sample the $x_{t-1}$ according to $x_{t}$ and $\hat{x}_{0}^{\varphi}$
$P(x_{t-1}|\hat{x}_{0}^{\varphi},x_{t})=\mathcal{N}(x_{t-1};\mu_{t}(\hat{x}_{0}^{\varphi},x_{t}),\tilde{\beta}_{t}\mathbf{I}).$ (6)
@@ 3.5 Diffusion Implicit Policy
We can observe that in Eq. 6, $x_{t-1}$ is sampled from $\mathcal{N}(\mu_{t},\tilde{\beta}_{t}\mathbf{I})$, thus we can adjust $\mu_{t}$ (i.e. the mean value of $x_{t-1}$) based on implicit policy and more suitable $x_{t-1}$ can be sampled accordingly. It is noteworthy, we need the final synthesized human motion $x_{0}$ to be consistent with the scene and achieve high reward. Thus, we propose to optimize $\mu_{t}$ through $\hat{x}_{0}^{\varphi}(\mu_{t},t-1,c)$ rather than $\mu_{t}$ itself. Here, we take $t-1$ as denoising step because $\mu_{t}$ is the mean value of the denoised $x_{t-1}$.
Thanks to the motion representation taken in Sec. 3.3, the reward functions are fully differentiable. Meanwhile, we find that optimizing $\hat{x}_{0}^{\varphi}(\mu_{t},t-1,c)$ shows better performance than directly modifying $\mu_{t}$ itself as previous work [46]. That is because direct optimization over $\mu_{t}$ does not ensure motion continuity. On the other side, optimizing $\mu_{t}$ via $\hat{x}_{0}^{\varphi}(\mu_{t},t-1,c)$ can help search a better distribution with higher interaction reward for the final synthesized motion $x_{0}$, and $\mu_{t}$ can be be adjusted as a whole in a GAN Inversion manner (given a Generator $x=G(z)$, adjust the latent code $z$ via loss $\mathcal{L}(x)$ for more desired output) [2, 3]. Similarly, given the diffusion model $\hat{x}_{0}^{\varphi}(\mu_{t},t-1,c)$, we optimize $\mu_{t}$ via $\mathcal{R}(\hat{x}_{0}^{\varphi})$ according the following formulation
@@
3.6 Multi-Task Motion Synthesis
@@
After a new round of motion synthesis, we obtain the generated motion $\{\hat{P}_{s}\}_{s=1:S}$. For more natural motion transition in the overlapping $H$ frames, we take a time-variant motion blending. Different from direct linear interpolation in the pose representation like priorMDM [36]. We only utilize linear interpolation for the human translation $\check{\tau}_{1:H}$. As for human orientation and joints’ rotations, we take the axis-angle representation $\bar{\theta}_{rot,s}=\tilde{\theta}_{\tilde{S}-H+s}$ and $\hat{\theta}_{rot,s}$ and convert them to the rotation matrix $\bar{M}_{rot,s}$ and $\hat{M}_{rot,s}$. The blending occur in the power space of rotation matrix which can be formulated as:

### v2

@@ Abstract
Scene-aware motion synthesis has been widely researched recently due to its numerous applications. Prevailing methods rely heavily on paired motion-scene data, while it is difficult to generalize to diverse scenes when trained only on a few specific ones. Thus, we propose a unified framework, termed Diffusion Implicit Policy (DIP), for scene-aware motion synthesis, where paired motion-scene data are no longer necessary. In this paper, we disentangle human-scene interaction from motion synthesis during training, and then introduce an interaction-based implicit policy into motion diffusion during inference. Synthesized motion can be derived through iterative diffusion denoising and implicit policy optimization, thus motion naturalness and interaction plausibility can be maintained simultaneously. For long-term motion synthesis, we introduce motion blending in joint rotation power space. The proposed method is evaluated on synthesized scenes with ShapeNet furniture, and real scenes from PROX and Replica. Results show that our framework presents better motion naturalness and interaction plausibility than cutting-edge methods. This also indicates the feasibility of utilizing the DIP for motion synthesis in more general tasks and versatile scenes. Code will be publicly available at https://github.com/jingyugong/DIP.
Introduction
@@
Figure 1: Policy learning frameworks. (a) Explicit policy is trained with paired motion-scene data. (b) Implicit policy optimizes the motion from random initialization. (c) Diffusion policy gradually denoise the motion. (d) Our diffusion implicit policy iteratively denoises and optimizes the motion to ensure motion naturalness, diversity, interaction plausibility without need for any paired motion-scene data.
@@
Synthesizing human motion in real 3D scenes has attracted significant attention in recent years (cao2020long; wang2021synthesizing; wang2022towards; zhao2023synthesizing; hu2024hoimotion), due to its wide applications in scene simulation, digital human animation, and virtual/augmented reality.
@@
Thanks to learning-based 3D perception (qi2017pointnet; zhao2021point), pioneers (cao2020long; wang2021synthesizing; zhao2023synthesizing) have attempted to synthesize motion in scenes with feasible human-scene interaction. However, in previous works, paired motion-scene data are required to learn scene-aware motion policies. The majority of prevailing methods (starke2019neural; zhao2023synthesizing) learn Explicit Policies to directly predict the desired motion based on current states and goals (Fig. 1 (a)). Some of them (wang2021synthesizing; wang2022towards) utilized a second-stage Implicit Policy optimization but sacrifice the motion naturalness for interaction plausibility (Fig. 1 (b)). Recent works (huang2023diffusion; wang2024move) utilized conditional Diffusion Policies to achieve better performance (Fig. 1 (c)), where massive paired motion-scene data is also necessary.
@@
In fact, human motion data (mahmood2019amass; lin2023motion; ren2023lidar) is far more abundant than paired motion-scene data (hassan2019resolving; wang2022humanise). Motion synthesis that relies heavily on paired data will inevitably suffer from limited diversity. Meanwhile, the generalization ability is hard to guarantee when trained on limited scenes and applied to various other scenes.
@@
Based on this observation, we propose a unified framework, termed Diffusion Implicit Policy (DIP) (Fig. 1 (d)), which disentangles human-scene interaction from motion synthesis during training and then integrate motion denoising with implicit policy optimization during inference. In this way, paired motion-scene data is no longer necessary, and motion naturalness and interaction plausibility can be ensured simultaneously for scene-aware motion synthesis.
@@
In the DIP, a motion diffusion model is employed to make the synthesized motion more and more natural throughout the entire denoising process(Fig. 2 (b)). We equipped the diffusion model with a ControlNet (zhang2023adding) branch to provide keyframe joint hints for historical motion and future goals. Following previous works (tevet2023human; xie2024omnicontrol), the diffusion model is designed to predict the original motion at each denoising step and then sample the denoised motion from a normal distribution accordingly. Thus, we can well utilize the stochastic process to pursue plausible human-scene interactions. Specifically, interaction-based reward functions are designed to assess the consistency between motions and scenes. These reward functions are used as implicit policy to optimize the sampling distribution, ensuring that the sampled denoised motion corresponds better to the 3D scenes at each denoising step. As the denoising process can also be treated as optimization for motion naturalness, the entire scene-aware motion synthesis can be framed as an optimization problem to pursue both motion naturalness and interaction plausibility.
@@
In addition, for long-term motion synthesis involving multiple tasks, we need to take historical motion as constrain when synthesizing future motion. To maintain continuity between historical and future motions, we employ a time-variant motion blending, where we interpolate the rotation matrix in the power space. Thus far, the proposed framework can synthesize long-term motion in general scenes without any training on paired motion-scene data.
For performance evaluation, we use scenes cluttered with furniture from ShapeNet (chang2015shapenet) to assess the ability on human-object interaction. We also take PROX (hassan2019resolving) and Replica (straub2019replica) to demonstrate the generalization ability in scene-aware motion synthesis. We compared the proposed method with prevailing works based on physical and perceptual scores. Comprehensive experiments support our claims and indicate that the synthesized motion produced by the proposed method demonstrates better performance.
@@
Related Work
Human Scene Interaction. Generating realistic and plausible human-scene interactions has been widely explored (savva2014scenegrok; savva2016pigraphs; zhang2020place; zhang2020generating; zhao2022compositional). PLACE (zhang2020place) modeled the proximity based on the distance between human body and 3D scene during interaction synthesis. POSA (hassan2021populating) designed a contact feature map for the human body, indicating the contact and semantic information for each vertex in the human mesh. COINS (zhao2022compositional) encoded the human body and 3D objects into a shared feature space and synthesizes diverse compositional interactions. Narrator (xuan2023narrator) modeled the correlation of 3D scene and text based on a scene graph for interaction generation.
@@
Motion Synthesis. Motion synthesis has been studied for a significant period (clavet2016motion; holden2017phase; starke2019neural; wang2019combining; guo2020action2motion; petrovich21actor; guo2022generating; wu2024doodle). TEMOS (petrovich2022temos) learned a shared latent space for motion and text alignment. MDM (tevet2023human) and MotionDiffuse (zhang2024motiondiffuse) introduced diffusion model into motion synthesis. Subsequent works (chen2023executing; xie2023towards; zhang2023remodiffuse; dai2024motionlcm; zhang2023finemogen; karunratanakul2023gmd; xie2024omnicontrol) further improved the controllability and quality of the generated results.
Thanks to the advancements in motion synthesis, we follow the MDM (tevet2023human) and extend it to scene-aware motion synthesis, using interaction-based reward functions as an implicit policy.
Scene-Aware Motion Synthesis. Synthesizing human motion in various scenes has garnered much attention recently (hassan2021stochastic; wang2022towards; zhang2022couch; hassan2023synthesizing; mullen2023placing; rempe2023trace; mir2024generating; cen2024generating; jiang2024scaling; yi2024generating; liu2024revisit). Wang $\it{et~al.}$ (wang2021synthesizing) utilized the PointNet (qi2017pointnet) to provide scene feature and optimized the entire motion after generation. GAMMA (zhang2022wanderings) and DIMOS (zhao2023synthesizing) learned a latent space for natural motion, and a policy network was trained over the latent motion space. SceneDiffuser (huang2023diffusion) proposed a scene-conditioned diffuser accompanied by a learning-based optimizer and planner. LAMA (lee2023locomotion) introduced a test-time optimization stage for controller network via reinforcement learning to predict the action cues for motion matching (clavet2016motion) and modification. AMDM (wang2024move) designed a two-stage framework with a scene affordance map as an intermediate representation.
@@
Method
Preliminary
Motion Representation. For human motion, we take the SMPL-X model (pavlakos2019expressive) to represent the pose at each frame. Here, we mainly consider the global orientation represented in axis-angle $\theta_{global}\in\mathbb{R}^{3}$, joint rotation in axis-angle $\theta_{j=1:21}\in\mathbb{R}^{63}$ and the translation $\tau\in\mathbb{R}^{3}$. Accordingly, for each frame $s$, the human pose can be defined as $P_{s}=\{\theta_{s,global},\theta_{s,j=1:21},\tau_{s}\}\in\mathbb{R}^{69}$, and the synthesized motion consisting of consecutive poses can be annotated as $\hat{P}=\{\hat{P}_{s}\}_{s=1:S}$. The body shape $\beta\in\mathbb{R}^{10}$ and hand pose $\theta_{h}\in\mathbb{R}^{24}$ are always keep the same as initial human body for simplicity. The first $K$ joints $J=J_{1:K}\in\mathbb{R}^{K\times 3}$ and body mesh with $V$ vertices $M(\tau,\theta_{global},\beta,\theta_{j},\theta_{h})\in\mathbb{R}^{V\times 3}$ are taken as auxiliary representation for human pose.
@@
where $\mathcal{N}$ is a normal distribution and $\alpha_{t=1:T}$ are hyper-parameters. The distribution of final noised motion $x_{T}$ will approximate to $\mathcal{N}(0,I)$. Like MDM (tevet2023human), we train a diffusion model to predict the original motion directly
@@
where $t$ is the time step and $a$ is the action label. For motion synthesis, the denoising procedure can be formulated as:
$P(x_{t-1}|\hat{x}_{0}^{\phi},x_{t})=\mathcal{N}(\mu_{t}(\hat{x}_{0}^{\phi},x_{t}),\tilde{\beta}_{t}\mathbf{I}),$ (3)
@@
Overview
Figure 2: (a) indicates the overall pipeline. Any feasible command will be decomposed into sub-tasks with action-object pairs. Then, we will synthesize future motion according to current sub-task. Last, the synthesized motions will be fused into the historical motion to obtain the final long-term motion. (b) presents the Diffusion Implicit Policy (DIP). In each iteration, the denoising step will make the synthesized motion appear more natural), and the implicit policy optimization will endow the motion with plausible interaction. The random sampling step can help the framework synthesize motion with diverse styles.
@@
We can first decompose the command into a list of sub-task (interaction behavior and object pair) via current LLMs (achiam2023gpt; touvron2023llama). For each sub-task, the human may go to some place or interact with the objects in the scene (e.g. , sitting on the chair).
@@
Given a sub-task, we will first locate the goal position using COINS (zhao2022compositional), and then fetch reward functions for implicit policy according to current action.
@@
We train a diffusion model conditioned on human action and keyframe joints to synthesize human motion with easy control. Later, we model the interaction-based reward functions and take them to optimize the sampling distribution of denoised motion at each denoising step. The motion prior from diffusion model and implicit policy from reward function are integrated together to synthesize scene-aware motion with desired interactions (as shown in Fig. 2 (b)).
@@
Given historical motion, the synthesized motion should be consistent with it. Thus, we design to derive the long-term motion via a motion blending where translation are interpolated linearly and rotation are blended in the matrix power space. By now, we can synthesize long-term motion in 3D scenes when all sub-tasks are completed.
Conditional Diffusion Model
@@
We take $x_{0}\in\mathbb{R}^{S\times 69}$ consisting of human translation, orientation, and joint rotations in $S$ frames as motion representation. To simplify the formulation of reward functions in our method, we train a diffusion model $\phi$ (Fig. 3 (a)) to predict the original motion $\hat{x}_{0}^{\phi}$ for each noised $x_{t}$ as shown in Eq. 2.
@@
For any original motion $x_{0}$, we take the local coordinate of motion to reduce representation redundancy. Specifically, we translate and horizontally rotate the motion based on the human pose in the first frame. In the transformed motion, the orientation of the human body in the first frame lie in the $yz$-plane ($y\geq 0$), and the pelvis is positioned at the origin. After motion synthesis, the motion in scenes can be derived via simple translation and horizontal rotation.
@@
Figure 3: Illustration of conditional diffusion model. A diffusion model is first trained conditioned on action, and then a ControlNet branch is taken to provide keyframe joints’ hint.
@@
To better maintain consistency with historical motion and achieve future goals, we also take a ControlNet branch (zhang2023adding) to provide the hint of keyframe human joints (available from motion data for training) that need to be controlled (Fig. 3 (b)). We choose to take the joint positions rather than joint rotation as the external hint for easier understanding of the space information. Concretely, we will first calculate the joints’ positions according to the original human motion, and then randomly select one from these joints in a few frames as the hint and others are padded with zeros, thus the input of ControlNet branch take the form of $\mathcal{J}=\{J_{s,k}\}_{s=1:S,k=1:K}$ where non-zero values provide the controlled joint position hints.
@@
For training the ControlNet branch, all its parameters are randomly initialized, and the link layers are initialized to zero to maintain the motion synthesis capability of the main branch. Meanwhile, all parameters in original motion diffusion model $\phi$ are loaded from a pre-trained model. The controlled diffusion model $\varphi$ is also supervised to predict the origin motion $x_{0}=\{P_{s}\}_{s=1:S}$. After fine-tuning, the controlled motion diffusion model can be formulated as
@@
For simplicity, the action $a$ and joint hint $\mathcal{J}$ are termed as the condition $c$. Later, the controlled diffusion model $\varphi$ are taken as motion prior to ensure motion naturalness.
@@
Reward Functions for Implicit Policy
Based on the aforementioned diffusion model, we introduce a set of differentiable reward functions that score the motions within scenes and serve as an implicit policy to enhance overall performance (as shown in Fig. 2 (b) blue part). Here, we primarily focus on three key aspects in the design of reward functions: motion continuity, goal achievement, and interaction plausibility.
To promote motion continuity, we introduce a Historical Motion Consistency reward $\mathcal{R}_{his}$ to ensure continuous transition between historical and synthesized motions, alongside a Smoothness reward to prevent abrupt changes with excessive acceleration $\mathcal{R}_{acc}$. Additionally, we implement a Goal Achievement reward $\mathcal{R}_{goal}$ that encourage the human to move closer to the target. We also incorporate a Contact reward $\mathcal{R}_{cont}$ and a Non-Penetration reward $\mathcal{R}_{pene}$ to facilitate plausible human-scene interactions, as well as a Non-skating reward $\mathcal{R}_{skt}$ to anchor the vertices of contact. The total reward function for implicit policy take the form of
$\begin{array}[]{rcl}\mathcal{R}_{ip}&\hskip-8.5359pt=&\hskip-8.5359pt\lambda_{his}\mathcal{R}_{his}+\lambda_{acc}\mathcal{R}_{acc}+\lambda_{goal}\mathcal{R}_{goal}\\ &&\hskip-8.5359pt+\lambda_{cont}\mathcal{R}_{cont}+\lambda_{pene}\mathcal{R}_{pene}+\lambda_{skt}\mathcal{R}_{skt}\end{array}$ (5)
@@
Diffusion Implicit Policy
By now, we utilize a diffusion model with ControlNet branch $\varphi$ to predict $\hat{x}_{0}^{\varphi}$. Thus, similar to Eq. 3, we can sample the $x_{t-1}$ according to $x_{t}$ and $\hat{x}_{0}^{\varphi}$
$P(x_{t-1}|\hat{x}_{0}^{\varphi},x_{t})=\mathcal{N}(\mu_{t}(\hat{x}_{0}^{\varphi},x_{t}),\tilde{\beta}_{t}\mathbf{I}).$ (6)
@@
In Eq. 6, $x_{t-1}$ is sampled from $\mathcal{N}(\mu_{t},\tilde{\beta}_{t}\mathbf{I})$, thus we can adjust $\mu_{t}$ (i.e. the mean value of $x_{t-1}$) based on implicit policy and more suitable $x_{t-1}$ can be sampled accordingly. It is noteworthy, we need the final synthesized human motion $x_{0}$ to be consistent with the scene and achieve high reward. Thus, we propose to optimize $\mu_{t}$ through $\hat{x}_{0}^{\varphi}(\mu_{t},t-1,c)$ rather than $\mu_{t}$ itself. Here, we take t1t-1 as denoising step because $\mu_{t}$ is the mean value of the denoised $x_{t-1}$.
Thanks to the motion representation that we take, the reward functions are fully differentiable. Meanwhile, we find that optimizing $\hat{x}_{0}^{\varphi}(\mu_{t},t-1,c)$ shows better performance than directly modifying $\mu_{t}$ itself as previous work (xie2024omnicontrol). That is because direct optimization over $\mu_{t}$ does not ensure motion continuity. On the other side, optimizing $\mu_{t}$ via $\hat{x}_{0}^{\varphi}(\mu_{t},t-1,c)$ can help search a better distribution with higher interaction reward for the final synthesized motion $x_{0}$, and $\mu_{t}$ can be be adjusted as a whole in a GAN Inversion manner (given a Generator $x=G(z)$, adjust the latent code $z$ via loss $\mathcal{L}(x)$ for more desired output) (bau2019semantic; bau2019seeing). Similarly, given the diffusion model $\hat{x}_{0}^{\varphi}(\mu_{t},t-1,c)$, we optimize $\mu_{t}$ via $\mathcal{R}(\hat{x}_{0}^{\varphi})$ according the following formulation
@@
Multi-Task Motion Synthesis
@@
After a new round of motion synthesis, we obtain the generated motion $\{\hat{P}_{s}\}_{s=1:S}$. For more natural motion transition in the overlapping $H$ frames, we take a time-variant motion blending. Different from direct linear interpolation in the pose representation like priorMDM (shafir2024human). We only utilize linear interpolation for the human translation $\check{\tau}_{1:H}$. As for human orientation and joints’ rotations, we take the axis-angle representation $\bar{\theta}_{rot,s}=\tilde{\theta}_{\tilde{S}-H+s}$ and $\hat{\theta}_{rot,s}$ and convert them to the rotation matrix $\bar{M}_{rot,s}$ and $\hat{M}_{rot,s}$. The blending occur in the power space of rotation matrix which can be formulated as: